1)
If variable costs per unit decreased because of a decrease in utility rates, the break-even point would Oa. decrease Ob. increase Oc. remain the same Od. increase or decrease, depending upon the percentage increase in utility rates
2)
If sales are $400,000, variable costs are 80% of sales, and operating income is $40,000, what is the operating leverage? Oa. 0.0 Ob. 1.3 Oc. 7.5 Od. 2.0
3)
If fixed costs are $561,000 and the unit contribution margin is $8.00, what is the break-even point in units if variable costs are decreased by $0.50 a unit? Oa. 70,125 Ob. 66,000 Ос. 60,000 Od. 74,800
4)
Strait Co. manufactures office furniture. During the most productive month of the year, 3,200 desks were manufactured at a total cost of $83,900. In the month of lowest production the company made 1,170 desks at a cost of $64,100. Using the high-low method of cost estimation, total fixed costs are Oa. $52,700 Ob. $64,100 Oc. $19,800 Od. $83,900
5)
Which of the following statements is true regarding fixed and variable costs? Oa. Both costs are constant when considered on a per-unit basis. Ob. Variable costs are constant in total, and fixed costs vary in total. Oc. Fixed costs are constant in total, and variable costs are constant per unit. Od. Both costs are constant when considered on a total basis
Answer #1

Solution 1:

If variable cost per unit decreased because of decrease in utility rates, the break even point would decrease.

Hence option a is correct.

Solution 2:

Operating leverage = Contribution margin / Operating income

= ($400,000*20%) / $40,000 = 2

Hence option d is correct.

Solution 3:

If variable cost are decreased by $0.50 per unit, new contribution margin per unit = $8 + $0.50 = $8.50 per unit

Fixed costs = $561,000

Break even units = Fixed costs / contribution margin per unit = $561,000 / $8.50 = 66000 units

Hence option b is correct.

Solution 4:

Variable cost per desk using high low method = (Cost at high level - Cost at low level) / (High level activity - Low level activity)

= ($83,900 - $64,100) / (3200 - 1170) = $9.75 per unit

Fixed cost = $83,900 - 3200*$9.75 = $52,700

Hence option a is correct.

Solution 5:

The true statement regarding fixed and variable cost is "Fixed costs are constant in total and variable costs are constant per unit"

Hence option c is correct.